OpenStudy (anonymous):

Multiply: 3x+12/8m• m^2/2x+8 3m/16 , m ≠ 0 3m/16 , m ≠ 0 and x ≠ −4 m(x+3)/8 , m ≠ 0 m(x+3)/8 , m ≠ 0 and x ≠ −4

OpenStudy (anonymous):

so what cancels?

OpenStudy (anonymous):

3x+12=3(x+4) 2x+8=2(x+4)

OpenStudy (anonymous):

3(x+4) m^2 -------- * ------------ 8 m 2(x+4)

OpenStudy (anonymous):

what?^